# Counterfactual Effect Generalization: A Combinatorial Definition

### v4

@@ 7 Discussion and Conclusion
The central methodological challenge in the Sciences, Policy-making and Design remains the evaluation of counterfactual statements (Did this treatment caused the result of interest? Did this policy?) The counterfactual definition of effects formulated sample properties necessary for effects to be free of selection biases, Eq.(1). The counterfactual definition of effects, Eq.(1), is based on the non-variation of effect backgrounds. The strategy in this work has been, in contrast, to focus on conditions where the amount of unobserved variation is guaranteed to be large (EV-increasing) and shared among factors (CF non-increasing). This alternative definition, as effects that remain invariant in the face of large external variation, can lead, instead, to generalizeable effect estimates. Under these conditions, causes can also be separated from confounders using effect variances, Principle.1. These considerations led to the notion of effect backgrounds, and the conditions under which they become sufficiently randomized. The perspective also allowed us to reconsider problems that have proven difficult for the pairwise counterfactual formulation, such as their sample size requirements and unobserved variable biases.
@@
To that end, we formulated a new Combinatorial and Modular Arithmetic model to connect sample incompleteness to loss of estimation performance. We formulated sample properties necessary for effect External Validity, Eq.(3), and their required sample sizes. They were used to define EV and variance statistics for individual counterfactual effect observations, as well as for sets of counterfactual observations (samples). Several concepts, only loosely justified in mainstream ML, such the as the hyperbolic functions  [Ribeiro, 2022a, LeCun et al., ] of neural networks, transformation group invariance  [Bietti and Mairal, 2019, Cohen and Welling, 2016], and 80-20 train-test sample split appears naturally in this purely combinatorial perspective. Cyclic permutations, and treatment-background ’convolutions’, Eq.(10), suggest a bridge between combinatorial and continuous multi-scale or frequency-based representations. The latter consists of one of the theoretical underpinning of neural networks and several other methods. This suggests a possible direction for the long-sought connection between causality and mainstream ML techniques  [Scholkopf et al., 2021]. We focused on confounding, external validity, and selection bias. Other concepts from the causal inference literature could be similarly implemented.
@@
A merit of the proposed approach is thus bringing causal and predictive concepts closer. We demonstrated generalizability bounds for popular supervised predictors and black-box explainers [Burkart and Huber, 2021]. Supervised prediction, black-box explanation and effect estimation approaches should state, side-by-side with their estimates, the conditions on which they are expected to hold (sample sizes, correlations, completeness, etc.). This is unknown for most current causal effect estimation solutions, IME explainers, and most out-of-sample prediction solutions. Until these are stated, it is difficult to, in practice, trust their outputs in real-life conditions. The approach allows us to model effects in both linear and non-linear data without the, rarely know and difficult to infer, parametric models for outcomes or strict independence among factors. The approach is related to the widely-employed Shapley value  [Burkart and Huber, 2021, Ribeiro, 2022a] and known U-Statistics  [Hoeffding, 1948, Yamato and Maesono, 1986], specialized for effect observations (not regressions), and generalized to non-i.i.d and partially observed data. We discussed out-of-sample, counterfactual, correlated and omitted variables prediction, as well as causal effect estimation, in simulations and an important real-world example. Correlated and counterfactual prediction observed, in particular, significant accuracy gains. A central result demonstrated here is that the EV of interventions are, to some extent, predictable from combinatorial properties of the populations they act upon. To a broad audience, Machine Learning methods are inherently limited due to their lack of generalizability and selection bias guarantees, when compared to, for example, experimental hypothesis testing. We believe the work could help build further connections between predictive and causal techniques.

### v5

@@ 8 Discussion and Conclusion
The central methodological challenge in the Sciences, Policy-making and Design remains the evaluation of counterfactual statements (Did this treatment caused the result of interest? Did this policy?) The counterfactual definition of effects  [Rubin, 2005, Morgan and Winship, 2007] formulate sample properties necessary for effects to be free of selection biases, Eq.(1). It is based on the fixation of sample effect backgrounds. The strategy in this work has been, in contrast, to focus on conditions where the amount of unobserved variation is guaranteed to be large (sampling is EV-increasing) and shared among factors (CF non-increasing). This alternative definition, as effects that remain invariant in the face of large external variation, can lead, instead, to generalizeable effect estimates. Under these conditions, causes can also be separated from confounders using effect variance estimates, Principle (1). These considerations led to the notion of effect background randomization, and sample conditions under which effects become sufficiently randomized. The perspective also allowed us to reconsider problems that have proven difficult for the standard counterfactual formulation, such as their sample size requirements and unobserved variable biases.
@@
The model allowed us to relate sample incompleteness to loss of estimation performance. We formulated sample properties necessary for effect External Validity, Eq.(3), and their required sample sizes. They were used to define EV and variance statistics for individual effect observations, as well as for sets of effect observations (samples). Several concepts, only loosely justified in mainstream ML, such the as the hyperbolic functions  [Ribeiro, 2022a, LeCun et al., ] of neural networks, transformation group invariance  [Bietti and Mairal, 2019, Cohen and Welling, 2016], and 80-20 train-test sample split appears naturally in this purely combinatorial perspective. Cyclic permutations, and treatment-background ’convolutions’, Eq.(10), suggest a bridge between combinatorial and continuous multi-scale or frequency-based representations. The latter consists of one of the theoretical underpinning of neural networks and several other methods. This suggests a possible direction for the long-sought connection between causality and mainstream ML techniques  [Scholkopf et al., 2021]. We focused on confounding, external validity, and selection bias. Other concepts from the causal inference literature could be similarly implemented.
@@
An advantage of the proposed approach is thus bringing causal and predictive concepts closer. We demonstrated generalizability bounds for popular supervised predictors and black-box explainers [Burkart and Huber, 2021]. Supervised prediction, black-box explanation and effect estimation approaches should state, side-by-side with their estimates, the conditions on which they are expected to hold (sample sizes, correlations, completeness, etc.). This is unknown for most current causal effect estimation solutions, IME explainers, and most out-of-sample prediction solutions. Until these are stated, it is difficult to, in practice, trust their outputs. The approach allows us to model effects in both linear and non-linear data without the, rarely know and difficult to infer, parametric models for outcomes or strict independence among factors. We discussed out-of-sample, counterfactual, correlated and omitted variables prediction, as well as causal effect estimation, in simulations and an important real-world example. Correlated and counterfactual prediction observed, in particular, significant accuracy gains. A central result demonstrated here is that the EV of interventions are, to some extent, predictable from combinatorial properties of the populations they act upon. To a broad audience, Machine Learning methods are inherently limited due to their lack of generalizability and selection bias guarantees, when compared to, for example, experimental hypothesis testing. We believe the work could help build further connections between predictive and causal techniques.